# TabSketchFM: Sketch-based Tabular Representation Learning for Data Discovery over Data Lakes

### v1

@@ 6. Experiments
    How does sketch-based TabSketchFM compare to existing cell-based models for building cross encoders for the tasks in LakeBench?
@@
    How can one use the cross encoders built from TabSketchFM for search over data lakes?
Table 2. Performance (avg $\pm$ stdev) of TabSketchFM vs. other baseline models on LakeBench averaged across five random seeds. Best performance is highlighted in bold, and the second best is underlined.
@@
We pretrain TabSketchFM using $4$ A100 $40$GB GPUs for $2$ days, until the model converged. Our pretrained model contains $118M$ parameters, similar to other row-based models in our experiments. We use patience of $5$, i.e., we consider the model as having converged if the validation loss does not decrease for more than $5$ epochs. As studied in the literature (Zhou et al., 2020), this not only reduces training time but also helps us avoid overfitting. For finetuning the cross encoders, for all tasks, we use $1$ A100 $40$GB GPU. Most finetuning tasks converged within $6$ hours using the same 5 epoch patience as pretraining except for the Wiki Union benchmark which took 24 hours. We use cross-entropy loss for classification tasks, mean squared error for regression tasks, and binary cross-entropy with logit loss for multi-class classification tasks. For each finetuning task, we create a cross-encoder (Du et al., 2022), which evaluates if a pair of tables can be joined, unioned or are subsets of each other.
6.1. TabSketchFM on LakeBench Tasks
@@ 6.1.1. Baselines
  • TaBERT provides two types of embeddings for a table: context embeddings corresponding to the tokens in the context, and column embeddings corresponding to each column in the table (Yin et al., 2020). For each table, we compute both embeddings for top $10,000$ rows. In Yin et al. (2020), a content snapshot approach is described which creates synthetic rows based on the n-gram matches of the input question with the cell values. We could not experiment with this feature, since 1. this technique is not part of the authors’ provided source code [^7] https://github.com/facebookresearch/TaBERT and 2. LakeBench tasks do not contain a question. So, we compute these embeddings for each table, mean-pooled over the context tokens and tabular columns respectively. The final embedding for the table pair is computed as the concatenation of context and column embeddings obtained for each table.
@@
  • TAPAS requires a natural language query and a table as input (Herzig et al., 2020). For our dataset discovery tasks, we sent an empty string as a natural language query and use $512$ serialized row based version of the table as the embedding.
  • TABBIE provides row and column embeddings of the table (Iida et al., 2021). For our work, we obtained the row embeddings for each row in the table. Following the original work, we use the first 30 rows and 20 columns of the table. These row embeddings are combined using the mean operation and the resulting vector is frozen as table embedding.
@@ 6.1.2. Results
Table 2 compares the performance of TabSketchFM with baseline models on LakeBench. For regression tasks, we report R2 statistics, and for (binary and multiclass) classification tasks, we report a weighted F1 score to handle skew in classes [^8] Our metrics were implemented from scikit-learn: https://scikit-learn.org/stable/modules/model_evaluation.html. The best performance is highlighted in bold and the second best is underlined. All the results presented are aggregated over 5 random seeds. We see that on five of the eight tasks, TabSketchFM outperformed all baselines, performed as well as the best methods on one, and performed second best on the remaining two.
@@
The binary classification adaptation of the TUS-SANTOS benchmark, was too easy and could be solved on the basis of column headers alone [^9] We added column headers for this benchmark, unlike in the literature, but removing headers made little difference to the results of tabular models., as evident from the performance of the Vanilla BERT model. Nevertheless, TAPAS and TABBIE perform relatively poorly on it, pointing to the importance of including the Vanilla BERT as a baseline. On the CKAN Subset benchmark, we find that the performance of most models was comparable to random guessing since the column headers were exactly the same, and most systems (except TaBERT) did not have a view of the entire dataset. On Wiki Union and on Spider-OpenData, TaBERT outperformed TabSketchFM. Even so, TabSketchFM achieved the second best performance on these two tasks.
We also performed an error analysis on table pairs that TabSketchFM got wrong on Wiki-Union, Spider open data and CKAN. In all these cases, mean hamming distances of MinHash sketches for positive pairs were equivalent to those from negative pairs; suggesting that the sketches alone did not contain sufficient information to discriminate those examples. TaBERT’s superior performance on the Wiki Union case, combined with the fact that lack of discrimination of MinHashes provides a clue to at least one weakness in the current version of TabSketchFM, which is that we do not consider any values from the column at all. As shown in Figure 4, two columns may be unionable when their values map to the same domain (e.g., different municipalities in Slovakia), with minimal overlap in actual values. In such cases, a system such as TaBERT can still determine that the cell values are similar, and belong to the same semantic type, whereas TabSketchFM cannot. Incorporating cell values is something we plan to address in future work.
In summary, the answer to  Q1 is that sketches-based TabSketchFM is in general superior to value-based methods for supervised data discovery in LakeBench. Note that this evaluation does not directly speak to performance on search tasks in data discovery; but it does evaluate the usefulness of sketches for building a good cross encoder. We turn next to investigating how different sketch types help specific tasks.
6.2. Effectiveness of sketch types on task types
To explore the relevance of the sketches and content snapshots for the different data discovery tasks, we conduct an ablation study where we use only one type of sketch. The results on LakeBench tasks with the random seed 0 are shown in Table 3. In  Table 4 we present another ablation study where we remove one sketch and retained the rest. Not surprisingly, MinHash sketches are crucial for join tasks, where MinHash alone seems to dominate overall performance of TabSketchFM. Subset selection clearly relies on the overlap of the data distribution; numerical sketches alone match the performance of TabSketchFM. Given that almost 69% of the columns were numerical in the subset benchmark, this outcome might be expected. The effect of content snapshot was less clear. Looking at Table 4, it is evident that the removal of content snapshot does in fact contribute somewhat to the Wiki join tasks. Numerical sketches contribute to ECB union which is perhaps because approximately 50% of the columns are not strings. Removal of MinHashes mirrors the result from earlier, which is, join tasks are most affected. Surprisingly, the removal of MinHashes affected the subset task; even though numerical sketches by themselves produced good performance. It is possible that there are complex interactions between the sketches. In summary, we answer Q2 from our ablation studies that different sketches play different roles across tasks, and may even interact to determining data discovery performance.
@@ 6.2. Effectiveness of sketch types on task types
Table 3. Importance of different sketches on TabSketchFM performance for the different tasks. Performance of sketches that are equivalent (within 2 points) to overall performance is highlighted in green. TUS-SANTOS is not considered because it can be performed based on column headers alone.
@@
Table 4. Removing a sketch to study the impact of it on TabSketchFM performance for the different tasks. Performance of sketches that declined more than 2 points compared to overall performance is highlighted in red.
@@
Further, in the technical report  [^10] Made available with other artifacts at https://anonymous.4open.science/r/tabsketchfm-6CDC/README.md we present a few more analysis to understand the robustness and sensitivity of TabSketchFM. Due to the lack of space, we only summarize a few important results here. First, we compare different hashing functions including: MMH3, FarmHash, SHA, Blake2S, xxhash and find that TabSketchFM is robust to the hashing implementations. Second, Pretraining is crucial to learn the data discovery tasks on smaller datasets. Third, large transformer-based models are crucial with sketches. Upon replacing the BERT model with MLP the performance is significantly impacted. Fourth, Minhash sketches support extremely large tables of enterprise data lakes. Using text tokens, shinglings, or Hashing Vectorizers is not scalable.
Figure 8. F1 on wiki join search for varying k.
Figure 9. F1 on Santos search for varying k.
Figure 10. F1 on TUS search for varying k.
6.3. Application to search
In the different search benchmarks we describe below, we combined the embeddings of columns derived from TabSketchFM models along with value embeddings for each column because as we describe in the introduction, using numerical and MinHash vectors as inputs to the transformer model made it difficult to include any cell values of a given column. To add cell values, which can contain important semantic information about a column, we created a single sentence by concatenating the top 100 values of a columnn as a sentence, and then embedded it using an off the shelf sentence embedding model (sentence-transformers/all-MiniLM-L12-v2). The embedding from TabSketchFM was concatenated to the value embeddings from the sentence embedder after normalizing them so the means and variances of the two vectors were in the same scale.
6.3.1. Join Search.
@@ 6.3. Application to search
Here, the task is to find all the tables from the data lake that can be joined on one or more columns with the given query table. Note that prior join works (Zhu et al., 2019; Esmailoghli et al., 2022a; Dong et al., 2021) evaluate the efficiency of searching for data lake columns having value overlap with the query column(s), and assume that the columns with overlapping values are always joinable. However, we are interested in whether it makes sense to join tables. For example, two columns: (i) explaining people’s Ages with values such as (50, 24, 66, 78, 92, 95) and (ii) explaining students’ marks out of 100 (rounded to integer) with values such as (78, 92, 95, 100, 20) can have overlapping values but it is not sensible to join them because they have different semantics. Because TabSketchFM is built to generate contextual embeddings of columns that include value overlap, column headings, its context with other columns, and so on, it should be possible to get rid of such useless joins. To our knowledge, there are no such benchmarks with annotated gold standards that consider sensibility. So, we construct a search benchmark named Wiki Join from Wikidata, by following the cell-entity mapping approach used in the curation of the Wiki Jaccard dataset (see Section 5.2). Precisely, we generate the tables and then create a list of all the pairs of joinable columns, i.e., columns with the same entity annotations and overlapping values. We consider the columns having the same entity annotations to be sensible to join. We assign each pair an associated overlap score, which is the Jaccard score of the entity annotation sets. For each column in the list of pairs, we create a ranked list of joinable columns in descending order of their scores.
We used existing join search methods such as LSHForest and JOSIE (Zhu et al., 2019), along with embeddings derived from fine tuned versions of other pretrained models such as TABERT as our baselines. A recent embedding-based join search system, WarpGate (Cong et al., 2023), is not available publicly. So, we implemented a similar Glove-embedding-based baseline (Pennington et al., 2014) (EmbedJoin) that embeds columns into an embedding space and performs a search over HNSW index (Malkov and Yashunin, 2020). Other openly available systems, such as MATE (Esmailoghli et al., 2022b) were not easy to install due to software versioning requirements that we could not resolve. Figure 8 shows the significant advantage of TabSketchFM over all other baselines.
6.3.2. Union Search.
@@
As in previous works (Fan et al., 2023; Hu et al., 2023), we use two publicly available benchmark datasets: $\mathbf{TUS_{Small}}$ and $\mathbf{SANTOS_{Small}}$ from  (Nargesian et al., 2018; Khatiwada et al., 2023a) because the larger versions of the dataset are not labeled. As a baseline, we compare against state-of-the-art table union search techniques: $\mathbf{D^{3}L}$ (Bogatu et al., 2020), SANTOS (Khatiwada et al., 2023a), and Starmie (Fan et al., 2023). We report results on $\mathbf{D^{3}L}$, SANTOS, and Starmie using their publicly available code. We could not compare against $AutoTUS$ (Hu et al., 2023), another recent union search system, because its code was not public at the time of writing this paper. For each technique, we use default parameters suggested in the respective papers. The results for TabSketchFM compared to the baselines show that TabSketchFM matched the best system (Fan et al., 2023) on SANTOS, see Figure 9. For TUS, on smaller K, STARMIE was better as shown in Figure 10, due to better recall.
6.3.3. Generalization across tasks and domains.

### v3

@@ IV Experiments
    How does sketch-based TabSketchFM compare to existing pretrained tabular models when finetuned for relevant table identification tasks in LakeBench?
@@
    How well do these fine tuned TabSketchFM models perform on data discovery tasks?
  4. Q4.
    Can TabSketchFM fine tuned on a single task and domain, adapt and generalize to a new domain?
TABLE II: Performance (avg $\pm$ stdev) of TabSketchFM vs. other baseline models on LakeBench averaged across five random seeds. Best performance is highlighted in bold, and the second best is underlined.
@@
We pretrain TabSketchFM using $4$ A100 $40$GB GPUs for $2$ days, until the model converged. [^8] The source code is available on https://github.com/ibm/tabsketchfm Our pretrained model contains $118M$ parameters, similar to other row-based models in our experiments. We use patience of $5$, i.e., we consider the model as having converged if the validation loss does not decrease for more than $5$ epochs. As studied in the literature [58] this not only reduces training time but also helps us avoid overfitting. For finetuning the cross encoders, for all tasks, we use one A100 $40$GB GPU. Most finetuning tasks converged within $6$ hours using the same 5 epoch patience as pretraining except for the Wiki Union benchmark which took 24 hours. Our code is accessible at https://anonymous.4open.science/r/tabsketchfm.
IV-A Fine-tuning models on LakeBench Tasks
@@ IV-A1 Baselines
  • TaBERT provides two types of embeddings for a table: context embeddings corresponding to the tokens in the context, and column embeddings corresponding to each column in the table [11]. For each table, we compute both embeddings for top $10,000$ rows. In [11], a content snapshot approach is described which creates synthetic rows based on the n-gram matches of the input question with the cell values. We could not experiment with this feature, since 1. this technique is not part of the authors’ provided source code [^9] https://github.com/facebookresearch/TaBERT and 2. LakeBench tasks do not contain a question. So, we compute these embeddings for each table, mean-pooled over the context tokens and tabular columns respectively. The final embedding for the table pair is computed as the concatenation of context and column embeddings obtained for each table.
@@
  • TAPAS requires a natural language query and a table as input [12]. For our dataset discovery tasks, we sent an empty string as a natural language query and use $512$ serialized row based version of the table as the embedding.
  • TABBIE provides row and column embeddings of the table [9]. For our work, we obtained the row embeddings for each row in the table. Following the original work, we use the first 30 rows and 20 columns of the table. These row embeddings are combined using the mean operation and the resulting vector is frozen as table embedding.
@@ IV-A2 Results
Table II compares the performance of TabSketchFM with baseline models on LakeBench. For regression tasks, we report R2 statistics, and for (binary and multiclass) classification tasks, we report a weighted F1 score to handle skew in classes [^10] Our metrics were implemented from scikit-learn: https://scikit-learn.org/stable/modules/model_evaluation.html. The best performance is highlighted in bold and the second best is underlined. All the results presented are aggregated over 5 random seeds. We see that on five of the eight tasks, TabSketchFM outperformed all baselines, performed as well as the best methods on one, and performed second best on the remaining two.
@@
The binary classification adaptation of the TUS-SANTOS benchmark, was too easy and could be solved on the basis of column headers alone [^11] We added column headers for this benchmark, unlike in the literature, but removing headers made little difference to the results of tabular models., as evident from the performance of the Vanilla BERT model. Nevertheless, TAPAS and TABBIE perform relatively poorly on it, pointing to the importance of including the Vanilla BERT as a baseline.
@@
TABLE III: Importance of different sketches on TabSketchFM performance for the different tasks. Performance of sketches that are equivalent (within 2 points) to overall performance is highlighted in green. TUS-SANTOS is not considered because it can be performed based on column headers alone.
@@
TABLE IV: Removing a sketch to study the impact of it on TabSketchFM performance for the different tasks. Performance of sketches that declined more than 2 points compared to overall performance is highlighted in red.
@@
On the CKAN Subset benchmark, we find that the performance of most models was comparable to random guessing since the column headers were exactly the same, and most systems (except TaBERT) did not have a view of the entire dataset. On Wiki Union and on Spider-OpenData, TaBERT outperformed TabSketchFM. Even so, TabSketchFM achieved the second best performance on these two tasks.
We also performed an error analysis on table pairs that TabSketchFM got wrong on Wiki-Union, Spider open data and CKAN. In all these cases, mean hamming distances of MinHash sketches for positive pairs were equivalent to those from negative pairs; suggesting that the sketches alone did not contain sufficient information to discriminate those examples. TaBERT’s superior performance on the Wiki Union case, combined with the fact that lack of discrimination of MinHashes provides a clue to at least one weakness in the current version of TabSketchFM, which is that we do not consider any values from the column at all. As shown in Figure 5, two columns may be unionable when their values map to the same domain (e.g., different municipalities in Slovakia), with minimal overlap in actual values. In such cases, a system such as TaBERT can still determine that the cell values are similar, and belong to the same semantic type, whereas TabSketchFM cannot. We show how to incorporate cell values externally in the section on search which addresses this issue.
In summary, the answer to  Q1 is that sketches-based TabSketchFM is in general superior to other pre-trained models for data discovery in LakeBench. Note that this evaluation does not directly speak to performance on search tasks in data discovery; but it does evaluate the usefulness of sketches for building a good model tuned for data discovery. We turn next to investigating how different sketch types help specific tasks, before returning to the question of how well these fine tuned models perform on search.
IV-B Effectiveness of sketch types on task types
To explore the relevance of the sketches and content snapshots for the different data discovery tasks, we conduct an ablation study where we use only one type of sketch. The results on LakeBench tasks with the random seed 0 are shown in Table III. In  Table IV we present another ablation study where we remove one sketch and retained the rest. Not surprisingly, MinHash sketches are crucial for join tasks, where MinHash alone seems to dominate overall performance of TabSketchFM. Subset selection clearly relies on the overlap of the data distribution; numerical sketches alone match the performance of TabSketchFM. Given that almost 69% of the columns were numerical in the subset benchmark, this outcome might be expected. The effect of content snapshot was less clear. Looking at Table IV, it is evident that the removal of content snapshot does in fact contribute somewhat to the Wiki join tasks. Numerical sketches contribute to ECB union which is perhaps because approximately 50% of the columns are not strings. Removal of MinHashes mirrors the result from earlier, which is, join tasks are most affected. Surprisingly, the removal of MinHashes affected the subset task; even though numerical sketches by themselves produced good performance. It is possible that there are complex interactions between the sketches. In summary, we answer Q2 from our ablation studies that different sketches play different roles across tasks, and may even interact to determining data discovery performance.
IV-C Application to search
We evaluate the embeddings from our finetuned TabSketchFM for three data discovery tasks: join search, union search, and subset search.
(a) F1 on wiki join search for varying k.
(b) F1 on Santos search for varying k.
(c) F1 on TUS search for varying k.
(d) F1 on Eurostat subset search for varying k.
Figure 4: F1 plot for search on various join, union and subset datasets.
TABLE V: F1, Precision & Recall for Wiki-join search.
Baseline Mean F1 P@10 R@10
TaBERT-FT 5.88 0.43 0.04
LSH-Forest 10.48 0.8 0.08
Josie 19.56 0.99 0.12
DeepJoin 18.88 0.96 0.11
WarpGate 18.58 0.95 0.11
SBERT 83.67 0.95 0.89
TabSketchFM 89.09 0.97 0.94
TabSketchFM-SBERT 92.81 0.98 0.99
TABLE VI: F1, Precision & Recall for SANTOS union search.
Baseline Mean F1 P@10 R@10
TaBERT-FT 36.64 0.63 0.46
TUTA-FT 25.34 0.43 0.3
Starmie 54.08 0.97 0.72
D3L 26.44 0.54 0.4
SANTOS 50.36 0.89 0.67
SBERT 53.86 0.97 0.73
TabSketchFM 51.38 0.92 0.69
TabSketchFM-SBERT 54.09 0.97 0.73
TABLE VII: F1, Precision & Recall for TUS union search.
Baseline Mean F1 P@60 R@60
TaBERT-FT 26.66 0.90 0.30
TUTA-FT 27.27 0.89 0.31
Starmie 27.48 0.96 0.32
D3L 18.98 0.75 0.21
SANTOS 20.83 0.81 0.23
SBERT 31.13 0.99 0.36
TabSketchFM 30.43 0.97 0.35
TabSketchFM-SBERT 30.72 0.99 0.35
TABLE VIII: F1, Precision & Recall for Eurostat subset search.
Baseline Mean F1 P@10 R@10
TABERT-FT 4.03 0.05 0.05
TUTA-FT 9.82 0.13 0.12
SBERT 43.12 0.56 0.51
TabSketch 49.96 0.59 0.53
TabSketch-SBERT 47.54 0.58 0.52
IV-C1 Join Search
@@
Here, the task is to find all the tables from the data lake that can be joined on one or more columns with the given query table. Note that prior join works [5, 59, 28] evaluate the efficiency of searching for data lake columns having value overlap with the query column(s), and assume that the columns with overlapping values are always joinable. However, we are interested in whether it makes sense to join tables. For example, two columns: (i) explaining people’s Ages with values such as (50, 24, 66, 78, 92, 95) and (ii) explaining students’ marks out of 100 (rounded to integer) with values such as (78, 92, 95, 100, 20) can have overlapping values but it is not sensible to join them because they have different semantics. Because TabSketchFM is built to generate contextual embeddings of columns that include value overlap, column headings, its context with other columns, and so on, it should be possible to get rid of such useless joins.
Figure 5: Sample subsets of four tables from our Wikidata-based table union and join benchmark (top), along with ground truth labels and mappings (bottom)
To our knowledge, there are no such benchmarks with annotated gold standards that consider if it is sensible to join [^12] A new benchmark very recently got published on this topic but that was after we had built the benchmark [49]. So, we construct a search benchmark named Wiki Join from Wikidata (by following the cell-entity mapping approach used in the curation of the Wiki Jaccard dataset [17], see Table I for dataset characteristics). We first generated the tables and then used the ground truth cell-entity mapping to create a list of all the pairs of columns with the same entity annotations. For example, consider Figure 5. All entities in first column of table A and first column of table B are mapped to Meteorite where as entities in table C are mapped to Locations. So even when there is an overlap in cell values of all tables (Aleppo appears in all the tables), only table A and B are considered sensible to join. Once the list of sensible joinable column pairs is obtained, we assign each pair an overlap score, which is the Jaccard score of the entity annotation sets. The pairs with overlap score greater than $0.5$ were considered as joinable in our dataset for a total of 5,512 queries. This dataset contains contains $46,521$ tables and $5,512$ query columns.
We compare the join search performance of TabSketchFM with two traditional join search methods—LSHForest and Josie [5]—and three neural methods–WarpGate [30], DeepJoin [29], and TaBERT. WarpGate is a recent embedding based method that performs joinable column search using SimHash LSH in the embedding space. It uses FastText embedding pretrained on WebTables to generate embeddings of column values. DeepJoin is a similar embedding base approach that finetunes a pretrained language model and uses HNSW [60] for indexing and efficient search. They experiment with FastText, BERT, DistilBERT and MPNet. In our experiments we use the FastText embedding as it was found to be the best performing embedding when used without finetuning. In addition to column values, DeepJoin also uses column names, table names and column statistics (max, min and average character length) while embedding; much like our sketches. For TaBERT baseline, we use the column embeddings obtained from the TaBERT model which was finetuned on the Wiki-Containment dataset from LakeBench in Section IV-A (TaBERT-FT). We could not include TUTA as it does not provide column embeddings.
Recognizing the power of pretrained models trained on massive corpora of text, we added an off-the-shelf sentence encoder SBERT (all-MiniLM-L12-v2) as a powerful baseline. We include a very simple approach of concatenating the top $100$ unique values in a column into a single sentence and encoding it to produced a column embedding. As seen in figure 4(a) this simple baseline was quite effective compared to many traditional hand crafted systems for data discovery, as well as other neural systems. For finding joins (and subsets), value overlap is a requirement, and data sketches do a very efficient job of finding overlap in large sets, whereas encoding many thousands of values in a sentence is inefficient and likely ineffective. To examine if there is any gain in combining the embeddings of columns derived from TabSketchFM models along with value embeddings for each column using sentence encodings, we also concatenated the two embeddings after normalizing them so the means and variances of the two vectors were in the same scale (TabSketchFM-SBERT).
Figure 4(a) shows the significant advantage of TabSketchFM over all other baselines, and furthermore, adding column embeddings using SBERT improved performance for joins. Table IV-C shows the mean F1 scores for all methods. In this and other benchmarks, the best performing system and the runner up were better than all other methods, using statistical significance tests. Table IV-C presents the Mean F1, Precision @ 10, and Recall @10 values.
[…]
Figure 6: Ranking of nearest tables
IV-C2 Union Search
@@
We use two publicly available benchmark datasets: $\mathbf{TUS_{Small}}$ and $\mathbf{SANTOS_{Small}}$ [4, 6] because the larger versions of the dataset are not labeled. As a baseline, we compare against state-of-the-art table union search techniques: $\mathbf{D^{3}L}$ [25], SANTOS [6], and Starmie [16] as well as the SBERT baseline explained earlier. We could not compare against $AutoTUS$ [26], a recent union search system, because its code was not public at the time of writing this paper.
To determine unionability, table embeddings from TabSketchFM did not yield strong results; perhaps because the union benchmarks had been created by creating subsets of the columns in a table or rows. Following the state-of-the-art unionability approach Starmie[16], we used column embeddings of TabSketchFM to determine if two tables are unionable. But we used a simpler technique than the bipartite graph matching algorithm introduced by Starmie. Our approach is defined in Figure 6, and works as follows:
  1. 1.
    For each column $c$ in a given table $t$, KnnSearch returns the nearest $k*3$ columns in the dataset. Note that we try to get a lot more columns than $k$ simply to increase the number of candidate columns we consider at this step because multiple columns from a single table might match a given column.
  2. 2.
    For a column $c$, ColumnNearestTables returns each table near $c$ and the distance of its closest column.
  3. 3.
    For a table $t$, NearestTables gathers the near tables for each column in $t$.
  4. 4.
    We then choose the best, first using $\textsc{Rank}_{1}$ to select the tables with the largest number of matching columns, and then, for ones where that is equal, $\textsc{Rank}_{2}$ chooses the closest, i.e. the smallest sum of column distances.
For each technique, we use default parameters suggested in the respective papers. For TUTA we had to use table embeddings because that was what the model made available. For both the TUTA and TaBERT, we use the model that was fine tuned on TUS-SANTOS dataset from LakeBench in Sec. IV-A(denoted by TUTA-FT and TaBERT-FT). The results for TabSketchFM compared to the baselines show that TabSketchFM matched the best system [16] on SANTOS, when we added embeddings from SBERT, see Figure 4(b), and Table IV-C for a summary across methods. For TUS, value embeddings from the off-the-shelf SBERT model performed very well, as did TabSketchFM when enhanced with value embeddings; see Figure 4(c) and Table IV-C. This suggests for union, column values alone are sufficient to find good unionable results; table specific embeddings do not seem to add much. Note that value embeddings out performed all the other systems, pointing to the power of existing pretrained models in embedding tabular data.
IV-C3 Subset Search
To our knowledge, there are no benchmarks for subset search. To ensure that the models we built generalize to search, we constructed a dataset for subset search, from Eurostat  [^13] https://data.europa.eu/en of 3,072 CSV files comprising the variety of data collected by the EU. These files serve as query tables for search. Data characteristics for this benchmark are provided in Table I. For each file, as shown in Figure 7, we created the following variants reflecting different types of subsets of the file, sampling 25%, 50% or 75% or all of the rows or columns in the table. When all rows or columns were kept, we shuffled either the rows or the columns (shown in purple). These last two variants were added to measure whether the embeddings capture table semantics, that is order invariance for rows and columns of a table. The 11 variants per file made up a search dataset of 38,904 files.
Figure 7: Sample table and generated subsets from Eurostat
(a) F1 on wiki join transfer on search for varying k.
(b) F1 on Santos transfer on search for varying k.
(c) F1 on TUS transfer on search for varying k.
(d) F1 on Eurostat transfer on search for varying k.
Figure 8: F1 plots for transfer across different K values. (a) WikiJoin (b) Santos (c) TUS (d) Eurostat
We follows the same procedure described in Figure 6 to find tables that were subsets of the query tables. Figure 4(d) and Table IV-C shows the results of the subset search on the fine tuned model trained on ckan-subset. As we see in the next section on transfer, this model turned out to be the weakest TabSketchFM models for the subset search problem, likely reflecting the quality of the dataset used for training The best performing TabSketchFM model was trained on union, on the European Central Bank data at 0.55. Still, we include it for consistency. In addition, the results suggest that for subsets, adding value embeddings of columns from SBERT actually made performance slightly worse. This highlights the importance of being able to customize embeddings for different data discovery tasks - the ability to add value embeddings to tabular embeddings seems to be useful for such customization.
One additional question we asked is whether the TabSketchFM models exhibited column order invariance, and row order invariance. Unsurprisingly, changing row order made no difference to TabSketchFM, all 3072 queried tables returned the row shuffled variants in their nearest neighbors set. In contrast, changing row order did influence SBERT embeddings; only 2809/3072 (91%) row shuffled variants were returned. Column order can matter to TabSketchFM models, since the encoder is a sequence model. For the column shuffled variants, 3059/3072 appeared in the nearest neighbors set (99.5%). The SBERT model of course did not consider the larger column context, and was hence unaffected by the shuffling of columns.
Overall, from our evaluations of $4$ data discovery datasets across traditional and neural approaches answers Q3, that embeddings from finetuned TabSketchFM can be used for data discovery tasks. As the embeddings are pretty small vectors the search in embedding space can be made extremely efficient; making it an ideal candidate for large data lakes.
IV-C4 Generalization across tasks and domains